Shift report 20010918 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Sep 18 2001 - 07:58:51 EDT

  • Next message: Flemming Videbaek: "BNL -- SECURITY ALERT UPDATE"

    ------------------------------------------------------------
    Date: Tue Sep 18 07:58:50 2001, Shift: 00:00-08:00
    Supervisor: J.I.Jordre, with: myself
    Log book#8 pages 151 - 152
    
    Shift Summary:
    ------------------------------------------------------------
    1. Store from last shift ended 02:37. Voted to end the store at 00:50 when ZDC rates were < 40Hz, i.e. ~1/4 of rate at beginning of the store. The last hour of the store MCR performed crystal collimator studies. Blue quench link interlock during dump. Thereafter recovery leading to MCR being unable to bring on RHIC RF accelerating cavity voltage. They had to call for an expert and are now setting up the machine again.
    2. Run 4780 (from last shift) with 303k trigger 1 events.
       Run 4781                   with  79k trigger 1 events.
    3. None other than the ones noted in previous shift.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Sep 18 2001 - 07:59:23 EDT